OpenStudy (anonymous):

Determine the slope and the y-intercept of the graph of the equation. -x + 10y = 40

OpenStudy (anonymous):

Remember the slope/intercept equation. y=mx+b 10y=-1x+40 10y/10=-1x/10+40/10 y=-x/10+4 I am a little rusty, I do not remember if m is the slope or if b is the slope. If you have reading material consult that, or you can check out purplemath.com, its a good math resource.